800 people are randomly chosen and 440 of them trust the president.

a. Find a point estimate of p

b. Find a 95% confidence interval to estimate the proportion of people who trust the president.

c. Conduct a hypothesis test to study whether the true proportion is greater than 0.5, α= 0.05

d. What is the p-value?
Answer #1

Solution - Given that, X = 440 n=800 7) P=* = 440 = 0.550 point estimate = 0.550 1- ² =1-0-550=0.450 is > At as o% confidencen=800 x=440 P = + = 440 - 0.55 Po=0.5 Hosp = 0.5 Haip > 0.5 This is right -tailed test Test statistic, z = o-po poll-po) - 0.